# Large-sample analysis of cost functionals for inference under the coalescent

### v1

@@ Abstract
The coalescent is a foundational model of latent genealogical trees under neutral evolution, but suffers from intractable sampling probabilities. Methods for approximating these sampling probabilities either introduce bias or fail to scale to large sample sizes. We show that a class of cost functionals of the coalescent with recurrent mutation and a finite number of alleles converge to tractable processes in the infinite-sample limit. A particular choice of costs yields insight about importance sampling methods, which are a classical tool for coalescent sampling probability approximation. These insights reveal that the behaviour of coalescent importance sampling algorithms differs markedly from standard sequential importance samplers, with or without resampling. We conduct a simulation study to verify that our asymptotics are accurate for algorithms with finite (and moderate) sample sizes. Our results also facilitate the a priori optimisation of computational resource allocation for coalescent sequential importance sampling. We do not observe the same behaviour for importance sampling methods under the infinite sites model of mutation, which is regarded as a good and more tractable approximation of finite alleles mutation in most respects.
@@ 1 Introduction
The coalescent (Kingman, 1982) is widely used in population genetics, either in its original form or in one of its numerous generalisations, to model or simulate the ancestral history (genealogy) of a sample of individuals. A crucial quantity for inference under the coalescent is the likelihood, or sampling probability, $p(\mathbf{n})$, i.e. the probability of observing a sample $\mathbf{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}$, with $n_{i}$ being the number of individuals carrying genetic type (allele) $i$, and $d$ being the number of possible alleles. Here we consider a finite number of alleles under recurrent mutation, and neglect other genetic forces such as selection and recombination. Even in this simple setting the sampling probability is not known explicitly, with the exception of so-called parent-independent mutation discussed in Remark 2.3 below. A recursive formula for $p(\mathbf{n})$ is available (Lundstrom et al., 1992; Sawyer et al., 1987), but unusable when the sample size $\|\mathbf{n}\|_{1}$ is even moderately large. Our interest is in the large-sample-size regime, to which we give precise meaning in Assumption 1.1.
@@
Because of the difficulty of computing the sampling probability exactly, even for moderate sample sizes, Monte Carlo methods have developed to estimate it. They broadly split into methods based on tree-valued Markov chain Monte Carlo, importance sampling and sequential Monte Carlo based on simulating coalescent trees sequentially from observed sequences at the leaves to the root, and approximate Bayesian computation which resorts to comparing observed and simulated summary statistics. Several review articles cover the range of methods available, and we direct the interested reader to Beaumont (2010); Marjoram and Tavaré (2006); Stephens (2007). We will develop an asymptotic description of a class of weighted functionals of the coalescent process, which admits analysis of importance sampling algorithms as a special case. Hence, we begin with an overview of coalescent importance sampling methods.
@@
The history of coalescent inference based on backward-in-time importance sampling starts with the Griffiths–Tavaré scheme (Griffiths and Tavaré, 1994b). Subsequently, Stephens and Donnelly (2000) developed a more efficient importance sampling algorithm by characterising the family of optimal but intractable proposal distributions, and by defining a tractable approximation. Their importance sampling scheme has since been extended in numerous ways, accounting for the infinite sites mutation model (Hobolth et al., 2008), selection (Stephens and Donnelly, 2003), recombination (Fearnhead and Donnelly, 2001; Griffiths et al., 2008), multiple mergers ($\Lambda$-coalescent) (Birkner et al., 2011; Koskela et al., 2015), and simultaneous multiple mergers ($\Xi$-coalescent) (Koskela et al., 2015).
@@
An entirely different approach to the approximation of the sampling probability consists of deriving series expansions amenable to asymptotics in regimes where some parameters are large. See for example (Jenkins and Song, 2009, 2010, 2012; Jenkins et al., 2015) for strong recombination, (Wakeley, 2008; Favero and Jenkins, 2023+; Fan and Wakeley, 2024) for strong selection, and (Wakeley and Sargsyan, 2009) for strong mutation. For the large-sample-size regime, the first order of the asymptotic expansion of the sampling probability is available (Favero and Hult, 2022) but it is expressed in terms of the generally unknown stationary density function of the Wright–Fisher diffusion. It does not seem possible to derive a more explicit expression, nor higher orders of the asymptotic expansion, by employing the classical techniques for the large parameters regimes mentioned above.
@@
Assumption 1.1 (Samples of large size).
@@
We consider samples of the form $n\mathbf{y}_{0}^{(n)}$, where $\mathbf{y}_{0}^{(n)}\in\frac{1}{n}\mathbb{N}^{d}$, and $n\in\mathbb{N}$ becomes large. We assume that the sequence $\mathbf{y}_{0}^{(n)}$ converges to some $\mathbf{y}_{0}\in\mathbb{R}_{+}^{d}$. For convenience we also assume $\|\mathbf{y}_{0}\|_{1}=1$. In this way, the size of the sample $n\mathbf{y}_{0}^{(n)}$, for large $n$, is approximately equal to $n$.
In this large-sample regime, we extend a previous convergence result on the block-counting process of the coalescent and the corresponding mutation-counting process (Favero and Hult, 2024) to include a sequence of costs. The convergence of general cost-weighted block-counting processes constitutes one of the two main results of this paper, Theorem 3.3. The proof is based on analysis of the tractable case of parent-independent mutation, and a change of measure between parent-independent and general recurrent mutation.
@@
We then use the cost framework we have developed to conduct a priori performance analysis of coalescent sequential importance sampling algorithms. The crucial idea for the analysis is based on the following interpretation. At each step, the discrepancy between a one-step proposal distribution and the intractable true sampling distribution can be viewed as the cost of that step. We write the sequential importance weights in terms of this cost sequence and employ our convergence result to study the asymptotic behaviour of the weights of classical importance sampling algorithms, particularly those of Griffiths and Tavaré (1994b) and Stephens and Donnelly (2000). This constitutes the second main theoretical contribution of the paper, Theorem 5.3.
@@
The idea of using a cost framework for the asymptotic analysis of importance sampling algorithms is inspired by the stochastic control approach to rare events simulation. This can be based on large deviations principles when the probability of the rare event is exponentially decaying, e.g. (Dupuis and Wang, 2004), or it can be based on Lyapunov methods in heavy-tailed settings, e.g. (Blanchet et al., 2012). These approaches are not applicable to the coalescent, necessitating the development of our bespoke approach based on convergence of cost functionals. While the main motivation for the construction of the cost framework is the analysis of importance sampling algorithms, the resulting limit of cost processes is generic and potentially of independent interest.
@@
Our theory makes the surprising prediction that, for large samples, normalised importance weights converge in distribution to 1 under mild conditions which both the Griffiths and Tavaré (1994b) and Stephens and Donnelly (2000) proposal distributions satisfy (c.f. Theorem 5.3 and Remark 5.4). Such convergence strongly suggests that the only contribution to overall importance weight variance arises from a relatively small number of sequential steps during which the number of remaining lineages in the coalescent tree is relatively small. This sets the coalescent apart from typical sequential importance sampling applications in which variance of importance weights grows exponentially in the number of steps (Doucet and Johansen, 2011). The fact that the behaviour of coalescent importance weights differs from standard settings has been observed before, and so-called stopping time resampling has been suggested as a remedy Chen et al. (2005); Jenkins (2012). Our results predict that the variance of coalescent importance weights remains non-standard even when stopping time resampling is employed.
@@
We conduct a simulation study to show that the predicted pattern of importance weight variance occurs in practice with moderate sample sizes. We make use of the effect by showing that coalescent sequential importance sampling methods can improved by using a small number of simulation replicates initially, and branching them out to a large number of replicates once the number of remaining extant lineages becomes small. The approach of targeting simulation replicates to those sequential steps which contribute to high variance is well-established (Lee and Whiteley, 2018), but typically relies on pilot runs to estimate one-step variances. Our theory facilitates its heuristic use for the coalescent without trial runs. In a similar vein, we show empirically that resampling, which typically reduces the growth of importance weight variance from exponential to linear in the number of steps (Doucet and Johansen, 2011), actually reduces the accuracy of the Stephens and Donnelly (2000) importance sampling algorithm.
@@
Finally, while our asymptotic theory is predicated on a finite number of alleles and recurrent mutation, we investigate whether similar empirical results hold for the so-called infinite sites model of mutation (see Section 6.2 for a description). The infinite sites model is regarded as a more tractable approximation of the finite alleles setting, but our results reveal a sharp difference between the two: state-of-the-art infinite sites importance sampling proposal distributions by Stephens and Donnelly (2000) and Hobolth et al. (2008) exhibit approximately exponential growth of importance weight variance with the number of sequential steps, resampling is effective at reducing Monte Carlo error, and non-uniform allocation of computational resources to different sequential steps does not improve performance. These results demonstrate that, from this perspective, the finite alleles and infinite sites models are not good approximations of each other. To carry out our infinite sites simulations, we derive some new computational complexity results for the proposal distribution of Hobolth et al. (2008) and show that pre-computing an explicit but large matrix reduces its complexity by an order of magnitude. The matrix in question is independent of observed data and can be reused across all simulations not exceeding a given sample size.
@@ 2.1 The coalescent and related sequences of interest
We consider the block-counting jump chain $\mathbf{H}=\{\mathbf{H}(k)\}_{k\in\mathbb{N}}\subset\mathbb{N}^{d}\setminus\{\bm{0}\}$ of the typed version of the coalescent, where $H_{i}(k)$ is the number of lineages of type $i$ after $k$ jumps in the ancestral history evolving backwards in time, and the coalescent is initialised from a starting configuration of types given by an observed sample $\mathbf{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}$, i.e. $\mathbf{H}(0)=\mathbf{n}$. The process stops when the most recent common ancestor (MRCA) of all individuals in the sample is reached at step
$\tau^{(n)}:=\inf\{k\in\mathbb{N}:\|\mathbf{H}(k)\|_{1}=1|\mathbf{H}(0)=\mathbf{n}\}.$
When not conditioning on $\mathbf{H}(0)$, the jump chain $\mathbf{H}$ has a tractable description as a forward-in-time process. It starts from one ancestor in the past with a type chosen from an initial type distribution, often the stationary distribution of the mutation matrix $P$, and evolves towards the present through mutation and branching events. The sampling probability $p(\mathbf{n})$ can be thought of as the probability that this forward process is in state $\mathbf{n}$ at the time of the first branching event which increases its number of lineages to $\|\mathbf{n}\|_{1}+1$. We record the forward and backward transition probabilities of the block-counting jump-chain $\mathbf{H}$ of the typed Kingman coalescent in Definition 2.1 and 2.2 below. See e.g. Stephens and Donnelly (2000); De Iorio and Griffiths (2004) for more details.
@@
$p(\mathbf{n}+\mathbf{v}|\mathbf{n})$ $=\mathbb{P}\left(\mathbf{H}(k)=\mathbf{n}+\mathbf{v}|\mathbf{H}(k+1)=\mathbf{n}\right)$ (2.1)
=$=\begin{cases}\frac{\|\mathbf{n}\|_{1}-1}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{j}}{\|\mathbf{n}\|_{1}}&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\quad j=1\dots d,\\ \frac{\theta}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{i}}{\|\mathbf{n}\|_{1}}P_{ij}&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\quad i,j=1\dots d,\\ 0&\text{ otherwise}.\end{cases}$
@@
$p(\mathbf{n}-\mathbf{v}|\mathbf{n})$ $=\mathbb{P}\left(\mathbf{H}(k+1)=\mathbf{n}-\mathbf{v}|\mathbf{H}(k)=\mathbf{n}\right)$ (2.2)
=$=\begin{cases}\frac{n_{j}(n_{j}-1)}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{1}{\pi[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\quad j=1\dots d,\\ \frac{\theta P_{ij}n_{j}}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{\pi[i|\mathbf{n}-\mathbf{e}_{j}]}{\pi[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\quad i,j=1\dots d,\\ 0,&\text{ otherwise},\end{cases}$
@@
Mutations are parent-independent when the type of the mutated offspring does not depend on the type of the parent, i.e. $P_{ij}=Q_{j},i,j=1,\dots,d$. In this special case, the sampling probability and the transition probabilities are explicitly known. In particular,
@@
The sequence of scaled block-counting Markov chains is defined as $\mathbf{Y}^{(n)}=\frac{1}{n}\mathbf{H}^{(n)}\subset\frac{1}{n}\mathbb{N}^{d},n\in\mathbb{N}$, where $n$ represent the sample size which we will take to grow to infinity.
@@
$M_{ij}^{(n)}(k)=\sum_{k^{\prime}=0}^{k-1}\mathbb{I}_{\{n\mathbf{Y}^{(n)}(k^{\prime})-n\mathbf{Y}^{(n)}(k^{\prime}+1)=\mathbf{e}_{j}-\mathbf{e}_{i}\}},$

### v2

@@ Abstract
The coalescent is a foundational model of latent genealogical trees under neutral evolution, but suffers from intractable sampling probabilities. Methods for approximating these sampling probabilities either introduce bias or fail to scale to large sample sizes. We show that a class of cost functionals of the coalescent with recurrent mutation and a finite number of alleles converge to tractable processes in the infinite-sample limit. A particular choice of costs yields insight about importance sampling methods, which are a classical tool for coalescent sampling probability approximation. These insights reveal that the behaviour of coalescent importance sampling algorithms differs markedly from standard sequential importance samplers, with or without resampling. We conduct a simulation study to verify that our asymptotics are accurate for algorithms with finite (and moderate) sample sizes. Our results constitute the first theoretical description of large-sample importance sampling algorithms for the coalescent, provide heuristics for the a priori optimisation of computational effort, and identify settings where resampling is harmful for algorithm performance. We observe strikingly different behaviour for importance sampling methods under the infinite sites model of mutation, which is regarded as a good and more tractable approximation of finite alleles mutation in most respects.
@@ 1 Introduction
The coalescent (Kingman, 1982) is widely used in population genetics, either in its original form or in one of its numerous generalisations, to model or simulate the ancestral history (genealogy) of a sample of individuals. A crucial quantity for inference under the coalescent is the sampling probability $p(\mathbf{n})$, i.e. the probability of observing a sample $\mathbf{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}$, with $n_{i}$ being the number of individuals carrying genetic type (allele) $i$, and $d$ being the number of possible alleles. In many statistical applications, the sampling probability constitutes the likelihood function for model parameters or other quantities of interest, such as the mutation rate or effective population size. Here we consider a finite number of alleles under recurrent mutation, and neglect other genetic forces such as selection and recombination. Even in this simple setting the sampling probability is not known explicitly, with the exception of so-called parent-independent mutation discussed in Remark 2.3 below. A recursive formula for $p(\mathbf{n})$ is available (Lundstrom et al., 1992; Sawyer et al., 1987), but intractable in practice when the sample size $\|\mathbf{n}\|_{1}$ is even moderately large. Our interest is in the large-sample-size regime, to which we give precise meaning in Assumption 1.1.
@@
Because of the difficulty of computing the sampling probability exactly, even for moderate sample sizes, Monte Carlo methods have been developed to estimate it. They broadly split into methods based on tree-valued Markov chain Monte Carlo, importance sampling and sequential Monte Carlo based on simulating coalescent trees sequentially from observed sequences at the leaves to the root, and approximate Bayesian computation which resorts to comparing observed and simulated summary statistics. Several review articles cover the range of methods available, and we direct the interested reader to Beaumont (2010); Marjoram and Tavaré (2006); Stephens (2007). We will develop an asymptotic description of a class of weighted functionals of the coalescent process, which admits analysis of importance sampling algorithms for large sample sizes as a special case. To our knowledge, this constitutes the first a priori analysis of coalescent importance sampling algorithms. We begin with an overview of coalescent importance sampling methods.
@@
The history of coalescent inference based on backward-in-time importance sampling starts with the Griffiths–Tavaré scheme (Griffiths and Tavaré, 1994b). Subsequently, Stephens and Donnelly (2000) developed a more efficient importance sampling algorithm by characterising the family of optimal but intractable proposal distributions, and by defining a tractable approximation. Their importance sampling scheme has since been extended in numerous ways, accounting for the infinite sites mutation model (Hobolth et al., 2008), selection (Stephens and Donnelly, 2003), recombination (Fearnhead and Donnelly, 2001; Griffiths et al., 2008), multiple mergers ($\Lambda$-coalescent) (Birkner and Blath, 2008; Birkner et al., 2011; Koskela et al., 2015), and simultaneous multiple mergers ($\Xi$-coalescent) (Koskela et al., 2015).
Importance sampling is closely connected to sequential Monte Carlo, which essentially amounts to generating many sequential importance sampling replicates in parallel and resampling them based on their importance weights to stochastically cull low-weight particles while duplicating high-weight ones (Chopin and Papaspiliopoulos, 2020, Chapter 9). Resampling is essential in many applications to prevent exponential growth of importance weight variance in the number of sequential steps (Doucet and Johansen, 2011). The behaviour of coalescent importance weights is known to differ from this regime (Fearnhead and Donnelly, 2001), and so-called stopping time resampling has been developed to make resampling applicable to coalescent models (Chen et al., 2005; Jenkins, 2012).
@@
An entirely different approach to the approximation of the sampling probability consists of deriving series expansions amenable to asymptotics in regimes where some parameters are large. See for example Jenkins and Song (2009, 2010, 2012); Jenkins et al. (2015) for strong recombination, Wakeley (2008); Favero and Jenkins (2025); Fan and Wakeley (2024) for strong selection, and Wakeley and Sargsyan (2009) for strong mutation. For the large-sample-size regime, the first order of the asymptotic expansion of the sampling probability is available (Favero and Hult, 2022) but it is expressed in terms of the generally unknown stationary density function of an associated Wright–Fisher diffusion. It does not seem possible to derive a more explicit expression, nor higher orders of the asymptotic expansion, by employing the classical techniques for the large parameters regimes mentioned above.
@@
Assumption 1.1 (Large-sample asymptotics).
We consider samples of the form $n\mathbf{y}_{0}^{(n)}$, where $\mathbf{y}_{0}^{(n)}\in\frac{1}{n}\mathbb{N}^{d}$, with $\frac{1}{n}\mathbb{N}^{d}$ being the lattice with spacing $\frac{1}{n}$, and we are interested in large-sample asymptotics, corresponding to $n\to\infty$. We assume that
$\lim_{n\to\infty}\mathbf{y}_{0}^{(n)}=\mathbf{y}_{0},$
@@
for some $\mathbf{y}_{0}\in\mathbb{R}_{+}^{d}$. For convenience we also assume $\|\mathbf{y}_{0}\|_{1}=1$. In this way, the size of the sample $n\mathbf{y}_{0}^{(n)}$, for large $n$, is approximately equal to $n$.
@@
In this large-sample regime, we extend a previous convergence result (Favero and Hult, 2024, Thm 2.1) on the block-counting process of the coalescent and the corresponding mutation-counting process to include a sequence of costs. The convergence of general cost-weighted block-counting processes constitutes one of the two main results of this paper, Theorem 3.3. Its proof is an extension and modification of that of (Favero and Hult, 2024, Thm 2.1), based on the analysis of the tractable parent-independent mutation case, followed by a change of measure between parent-independent and general recurrent mutation. This extended convergence result connects the asymptotics of Favero and Hult (2024) with practical coalescent sequential importance sampling algorithms, which we demonstrate by conducting an a priori analysis of their performance. To our knowledge, this work is the first tractable description of these algorithms outside the special case of parent-independent mutation. Previous work has focused on approximating the optimal but intractable proposal distribution Hobolth et al. (2008); Stephens and Donnelly (2000), whereas our focus is describing the variance arising from the use of tractable proposal distributions. The crucial idea for the analysis is based on the following interpretation. At each step, the discrepancy between a one-step proposal distribution and the intractable true sampling distribution can be viewed as the cost of that step. We write the sequential importance weights in terms of this cost sequence and employ our convergence result to study the asymptotic behaviour of the weights of classical importance sampling algorithms, particularly those of Griffiths and Tavaré (1994b) and Stephens and Donnelly (2000). This constitutes the second main theoretical contribution of the paper, Theorem 5.3.
@@
The idea of using a cost framework for the asymptotic analysis of importance sampling algorithms is inspired by the stochastic control approach to rare events simulation. This can be based on large deviations principles when the probability of the rare event is exponentially decaying, e.g. Dupuis and Wang (2004), or it can be based on Lyapunov methods in heavy-tailed settings, e.g. Blanchet et al. (2012). These approaches are not applicable to the coalescent, necessitating the development of our bespoke approach based on convergence of cost functionals. While the main motivation for the construction of the cost framework is the analysis of importance sampling algorithms, the resulting limit of cost processes is generic and potentially of independent interest.
@@
Our theory makes the surprising prediction that, for large samples, normalised importance weights converge in distribution to 1 under mild conditions which both the Griffiths and Tavaré (1994b) and Stephens and Donnelly (2000) proposal distributions satisfy (c.f. Theorem 5.3 and Remark 5.4). Such convergence strongly suggests that the only contribution to overall importance weight variance arises from a relatively small number of sequential steps during which the number of remaining lineages in the coalescent tree is small. This sets the coalescent apart from typical sequential importance sampling applications in which variance of importance weights grows exponentially in the number of steps (Doucet and Johansen, 2011). The fact that the behaviour of coalescent importance weights differs from standard settings has been observed before (Fearnhead, 2008), but our results predict that the variance of coalescent importance weights remains non-standard even when stopping time resampling is employed.
@@
We conduct a simulation study to show that the predicted pattern of importance weight variance occurs in practice with moderate sample sizes. We make use of the effect by showing that coalescent sequential importance sampling methods can be improved by using a small number of simulation replicates initially, and branching them out to a large number of replicates once the number of remaining extant lineages becomes small. The approach of targeting simulation replicates to those sequential steps which contribute to high variance is well-established (Lee and Whiteley, 2018), but typically relies on pilot runs to estimate one-step variances. Our theory facilitates its heuristic use for the coalescent without trial runs. In a similar vein, we show empirically that resampling, which typically reduces the growth of importance weight variance from exponential to linear in the number of sequential importance sampling steps (Doucet and Johansen, 2011), actually reduces the accuracy of the importance sampling algorithm of Stephens and Donnelly (2000), even when stopping time resampling is used. Thus, the practical contributions of our theory are (i) a heuristic for optimising the allocation of computational resources for finite alleles importance sampling, and (ii) a caution against employing resampling (including stopping time resampling) without a case-by-case assessment of its impact. The latter is significant because resampling is widely regarded as essential for good performance of practical importance sampling schemes, to which we present a clear and practical counterexample in Section 6.1.
@@
Finally, while our asymptotic theory is predicated on a finite number of alleles and recurrent mutation, we investigate whether similar empirical results hold for the so-called infinite sites model of mutation (see Section 6.2 for a description). The infinite sites model is a more tractable approximation of the finite alleles setting, but our results reveal a sharp difference between the two: state-of-the-art infinite sites importance sampling proposal distributions by Stephens and Donnelly (2000) and Hobolth et al. (2008) exhibit approximately exponential growth of importance weight variance with the number of sequential steps, resampling is effective at reducing Monte Carlo error, and non-uniform allocation of computational resources to different sequential steps does not improve performance. These results demonstrate that the finite alleles and infinite sites models do not resemble each other when from the perspective of importance sampling methods. To carry out our infinite sites simulations, we derive some new computational complexity results for the proposal distribution of Hobolth et al. (2008) and show that pre-computing an explicit but large matrix reduces its complexity by an order of magnitude. The matrix in question is independent of observed data and can be reused across all simulations not exceeding a given sample size.
@@ 2.1 The coalescent and related sequences of interest
For a fixed sample size $n\in\mathbb{N}$, we consider the block-counting jump chain $\mathbf{H}^{(n)}=\{\mathbf{H}^{(n)}(k)\}_{k\in\mathbb{N}}\subset\mathbb{N}^{d}\setminus\{\bm{0}\}$ of the typed version of the coalescent, where $H_{i}^{(n)}(k)$ is the number of lineages of type $i$ after $k$ jumps in the ancestral history evolving backwards in time, and the coalescent is initialised from a starting configuration of types given by an observed sample $\mathbf{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}$, i.e. $\mathbf{H}^{(n)}(0)=\mathbf{n}$. The process stops when the most recent common ancestor (MRCA) of all individuals in the sample is reached at step
$\tau^{(n)}:=\inf\{k\in\mathbb{N}:\|\mathbf{H}^{(n)}(k)\|_{1}=1|\mathbf{H}^{(n)}(0)=\mathbf{n}\}.$
When not conditioning on $\mathbf{H}^{(n)}(0)$, the jump chain $\mathbf{H}^{(n)}$ has a tractable description as a forward-in-time process. It starts from one ancestor in the past with a type chosen from an initial type distribution, often the stationary distribution of the mutation matrix $P$, and evolves towards the present through mutation and branching events. The sampling probability $p(\mathbf{n})$ can be thought of as the probability that this forward process is in state $\mathbf{n}$ at the time of the first branching event which increases its number of lineages to $\|\mathbf{n}\|_{1}+1$. We record the forward and backward transition probabilities of the block-counting jump-chain $\mathbf{H}^{(n)}$ of the typed Kingman coalescent in Definition 2.1 and 2.2 below. See e.g. Stephens and Donnelly (2000); De Iorio and Griffiths (2004) for more details.
@@
$p(\mathbf{n}+\mathbf{v}|\mathbf{n})$ $=\mathbb{P}\left(\mathbf{H}^{(n)}(k)=\mathbf{n}+\mathbf{v}|\mathbf{H}^{(n)}(k+1)=\mathbf{n}\right)$ (2.1)
[…]
where $\mathbf{e}_{i}$ is the canonical unit vector with a one at position $i$ and zeros elsewhere.
@@
$p(\mathbf{n}-\mathbf{v}|\mathbf{n})$ $=\mathbb{P}\left(\mathbf{H}^{(n)}(k+1)=\mathbf{n}-\mathbf{v}|\mathbf{H}^{(n)}(k)=\mathbf{n}\right)$ (2.2)
[…]
@@
Mutations are parent-independent when the type of the mutated offspring does not depend on the type of the parent, i.e. $P_{ij}=Q_{j},i,j\in\{1,\dots,d\}$. In this special case, the sampling probability and the transition probabilities are explicitly known. In particular,
@@
The sequence of scaled block-counting Markov chains is defined as $\mathbf{Y}^{(n)}=\frac{1}{n}\mathbf{H}^{(n)}\subset\frac{1}{n}\mathbb{N}^{d},n\in\mathbb{N}$, where $n$ represents the sample size which we will take to grow to infinity.
@@
$M_{ij}^{(n)}(k)=\sum_{k^{\prime}=0}^{k-1}\mathbb{I}(n\mathbf{Y}^{(n)}(k^{\prime})-n\mathbf{Y}^{(n)}(k^{\prime}+1)=\mathbf{e}_{j}-\mathbf{e}_{i}),$